OpenStudy (e.cociuba):

Solve the equation by completing the square: 2n^2 + 12n + 10=0

OpenStudy (anonymous):

n^2 + 6n + 5 = 0 (n+5) (n+1) = 0 n = -5 OR n = -1

OpenStudy (anonymous):

first we divided by 2

OpenStudy (anonymous):

to make it simple

OpenStudy (e.cociuba):

So is that a faster way to solve the prob?

OpenStudy (anonymous):

it depends on the problem

OpenStudy (e.cociuba):

KK thanks sooo much! :)

OpenStudy (anonymous):

you welcome :)

OpenStudy (e.cociuba):

:)